OpenStudy (anonymous):

9/32 minus 1/16 Help

OpenStudy (anonymous):

0.21875

OpenStudy (anonymous):

@kittypirate wonts a fraction!

OpenStudy (anonymous):

7/32

OpenStudy (anonymous):

@limkyle thanks so much

OpenStudy (anonymous):

9/32 - 1/16 = ? Get the LCD 9/32 - 2/32 = 7/32

OpenStudy (anonymous):

Thanks evryone